A cargo ship off the coast of japan is moving at 37.5km/h [217] with the north current moving at 9km/h [62.5].determine vector components?determine vector components for the current.whats the velocity of the boat in the end?


1
Expert's answer
2021-02-24T16:59:24-0500
"v_x=37.5\cos{217}+9\cos{62.5}=-25.79\frac{km}{h}\\v_y=37.5\sin{217}+9\sin{62.5}=-14.59\frac{km}{h}\\v=\sqrt{25.79^2+14.59^2}=29.6\frac{km}{h}"

"\theta=270-\arctan{\frac{14.59}{25.79}}=240.5\degree"